# Novel method for in-situ drift velocity measurement in large volume TPCs: the Geometry Reference Chamber of the NA61/SHINE experiment at CERN

### v1

@@ Abstract
This paper presents a novel method for low maintenance, low ambiguity in-situ drift velocity monitoring in large volume Time Projection Chambers (TPCs). The method was developed and deployed for the $40\,\mathrm{m}^{3}$ TPC tracker system of the NA61/SHINE experiment at CERN, which has a one meter of drift length. The method relies on a low-cost multi-wire proportional chamber (MWPC) placed downstream of the TPCs to be monitored. The drift velocity is then determined by matching the reconstructed tracks in the TPC to the hits of the pertinent monitoring chamber, called Geometry Reference Chamber (GRC), which is then used as a differential length scale. An important design requirement on the GRC was minimal added complexity to the existing system, in particular, compatibility with Front-End Electronics (FEE) cards already used to read out the TPCs. Moreover, the GRC system was designed to operate both in large and small particle flux. The system is capable of monitoring the evolution of the in-situ drift velocity down to a one permil precision, with a few minutes of time sampling.
@@ 1 Introduction
The NA61 experiment, also called the SPS Heavy Ion and Neutrino Experiment (SHINE) [2], is a large-acceptance hadron spectrometer experiment at the Super Proton Synchrotron (SPS) accelerator at CERN. Its physics program includes measurements for heavy-ion physics, as well as measurements on particle production in hadron-nucleus collisions, with an emphasis on the application of those measurements in flux predictions for long-baseline neutrino oscillation experiments (T2K, DUNE) and for cosmic ray observatories (Pierre Auger Observatory). These physics programs demand the operability of NA61/SHINE in both high and low particle multiplicity environments.
@@
The main components of the NA61/SHINE experiment are two large superconducting bending magnets and a system of large volume Time Projection Chambers (TPCs). As the drift dimension of the TPCs is considerably large (${\gtrsim}1\,$meter), in order to achieve necessary position resolution in the drift direction, the in-situ drift velocity inside the TPCs needs to be monitored needs to me monitored at the permil level. Since the drift velocity slowly evolves with time, its evolution needs to be followed with a few minutes of time sampling. In this paper, we describe a cost-efficient supplementary detector system specifically developed for this purpose.
@@
One of the default methods for monitoring drift velocity in a TPC is to analyze its exhaust gas with a small probe chamber [1, 2, 3, 4] and then predict the in-situ drift velocity of the TPC via the known pressure, temperature and drift field in the TPC gas. This method dates back to the large-volume TPCs at LEP. NA61/SHINE also uses this technique to monitor the drift velocity in its TPCs, but due to the limited absolute systematic accuracy of the method it is mainly used for monitoring the time stability of the gas composition. [^1] The so-called normalized drift velocity, estimated via exhaust analyzer, is very sensitive to changes in the working gas mixture composition. This quantity is the drift velocity extrapolated to normal pressure ($101325\,k\mathrm{Pa}$) and normal temperature ($20\,{}^{\circ}C$), and to a fixed drift field. The exhaust method is prone to uncertainties in pressure and temperature measurements in the large-volume chambers, and is also vulnerable to contaminations along the exhaust line and system outgassing.
@@
Due to the systematic uncertainties of the exhaust analyzer method, typically it is complemented by absolute in-situ drift velocity measurements. For this, typically a UV laser system is used: ionizing laser beam tracks are projected into the sensitive volume at known locations, and the drift velocity can then be measured from the apparent positions of these beams in terms of TPC drift time. Direct measurement methods using laser systems are common [5, 3] and very accurate. UV laser-based drift velocity and geometry monitoring systems, however, are rather hard to implement or upgrade on existing chambers retrospectively. [^2] Due to historical reasons, NA61/SHINE is not equipped with such a UV laser based system. Moreover, the cost and development time for such systems is substantial.
@@
Along with other CERN experiments, the NA61/SHINE facility was also significantly upgraded during the Long Shutdown 2 (LS2) period (2019-2021) of the accelerator complex. Since during the LS2 the subdetectors were moved, precise alignment calibration of the chambers was required, as was a system for accurately and reliably determining their drift velocities. As the TPCs were already constructed, minimally-invasive measurement systems were preferred. Our chose was a simple solution: a planar detector with fixed segmentation along the TPC drift direction, called the Geometry Reference Chamber (GRC), was to be placed downstream of the TPC system. TPC tracks reconstructed with an approximately known drift velocity were extrapolated to the GRC and paired with the GRC hits. If the drift velocity estimate differs from the true drift velocity inside the TPC, the mismatch of the drift coordinate of the hits becomes worse with increasing drift depth. The correction between the drift velocity estimate and the true drift velocity then can be calculated from the slope of this mismatch scatter plot. That is, the GRC is used as a differential length scale in order to convert the drift time, measured by the TPC readout, to drift depth. It is quite advantageous that the GRC method is differential: an accurate alignment of the TPC against the GRC is not needed. The concept is illustrated in Fig.1.333Once the drift velocity has been calibrated with such method in one of the TPC chambers in the detector complex, the other chambers can be calibrated against each-other, using the already calibrated chambers as geometry reference. Moreover, the remaining geometry calibration constants of the TPC system can also be determined, such as the trigger latency ($t_{0}$), alignment shift and alignment angles for each chamber, see Appendix A and B.
@@
Figure 1: (Color online) Illustration of the Geometry Reference Chamber (GRC) concept. The particle tracks in the TPC chamber are reconstructed with an initial estimate for the drift velocity and are paired to the hits of a segmented detector, the GRC. In case the drift velocity estimate differs from the true in-situ drift velocity inside the TPC, the mismatch along the drift coordinate gets gradually worse at increasing drift depths. The slope of the drift coordinate mismatch against the drift coordinate as measured by the GRC provides the correction between the true in-situ drift velocity versus the initial estimate.
@@
In order to implement the GRC concept, several design requirements were considered. To reduce the development time and cost of the readout electronics, a solution using the existing TPC readout Front-End Electronics (FEE) was chosen. In order to reduce the number of readout channels, we decided to implement a simple Multi-Wire Proportional Chamber (MWPC) with cartesian readout. This required two issues to be addressed. Firstly, the signal formation in the GRC had to be slowed down sufficiently such that the signal persists until the TPC electronics begins to sample the charges. This is not entirely trivial to achieve, since TPC readout electronics typically have a microsecond-level trigger latency with respect to the particle passage time, during which the signal may already disappear from the GRC. In order to mitigate this issue, a drift volume with adjustable drift field was introduced in the GRC design, transverse to the particle passage direction, and with that the GRC signal duration could be controlled. Secondly, the cartesian readout scheme tends to be problematic whenever the particle occupancy is large. In order to overcome this limitation, the GRC acceptance was designed to be variable: for low-multiplicity type of collisions, the entire GRC acceptance can be used, whereas for high-multiplicity collisions, individual anode wires (sense wires) of the GRCs can be disabled, down to a single amplifying wire. This helps reduce the number of combinatorial ghost hits.
@@
The structure of the paper is as follows. In Section 2 an overview of the NA61/SHINE experimental facility is given. In Section 3 the GRC design requirements, parameters, and implementation are discussed. In Section 4 the operational performance is shown. In Section 5 a conclusion is provided. The paper is closed by Appendix A and Appendix B discussing the details of systematic errors and the calibration of further geometry-related TPC calibration constants, where the GRC chamber plays a role.
@@ 2 The NA61/SHINE experimental facility
The NA61/SHINE [2] is a fixed-target hadron spectrometer experiment at the CERN SPS accelerator. Large parts of its tracking devices were inherited from a previous experiment called NA49 [3]. Its physics program covers the study of strongly-interacting matter via heavy-ion collisions and measurements of identified particle production spectra in hadron-nucleus collisions as reference data for flux prediction in long baseline neutrino experiments (T2K, DUNE) and large area cosmic ray observatories (Pierre Auger Observatory).
@@
An overview of the NA61/SHINE experiment is shown in Fig.2. Two large superconducting bending magnets (Vertex-I and II) are responsible for particle deflection for charge and momentum determination. Their total maximum bending power is ${\sim}9\,\mathrm{Tm}$ (up to $1.5\,\mathrm{Tesla}$ magnetic field in Vertex-I and $1.1\,\mathrm{Tesla}$ in Vertex-II). A target holder with target-in / target-out moving capability sits just upstream of the first Vertex TPC. Thin targets can be placed inside a silicon Vertex Detector (VD) for precise vertex determination. NA61/SHINE also has the ability to measure interactions in extended replica targets for long baseline neutrino experiments. The tracking devices for spectrometry are composed of eight large volume TPCs (total ${\sim}40\,\mathrm{m}^{3}$ and ${\sim}1\,\mathrm{m}$ drift length), capable of performing both tracking and $\mathrm{d}E/\mathrm{d}x$ measurements. An MRPC-based Time-of-Flight wall (L-ToF) provides further particle identification (PID) capabilities around mid-rapidity, whereas a scintillator based Time-of-Flight wall covers (F-ToF) covers the forward phase space, enabling two-dimensional (ToF+$\mathrm{d}E/\mathrm{d}x$) separation of particle species at large parts of the acceptance. A calorimeter is placed at the end of the beamline, called the Projectile Spectator Detector (PSD), which helps to characterize collision centrality in heavy-ion collisions, and consists of two compartments (MPSD and FPSD) for optimal shower containment. Upstream of the target position, a set of beam position detectors (BPDs), scintillator and Cherenkov detectors serve as beam and beam PID trigger (not shown in the figure). On the beamline between VTPC-1 and the GapTPC, a small plastic scintillator with a $1\,\mathrm{cm}$ diameter serves as an interaction trigger in most collision types (S4). In rare run settings, a very similar scintillator (S5) just upstream of MPSD is used instead. For heavy-ion runs the scintillator (S3) just downstream of VD is used for interaction definition through detecting loss of charge (i.e., by sensing a decreased $Z^{2}$ with respect to the beam nuclei). The VD, the new silicon BPDs, the MRPC-based L-ToF, the MPSD and FPSD were introduced during the Long Shutdown 2 (LS2) upgrade period between 2019-2021. Also during LS2, the data acquisition (DAQ) was upgraded, allowing for an event recording rate up to $1.8\,k\mathrm{Hz}$ for the detector complex.
@@
Figure 2: (Color online) The NA61/SHINE detector configuration, after the Long Shutdown 2 (LS2) upgrade, i.e. since 2022. The drawing shows the top view of the bending plane. That is, the approximately homogeneous magnetic field inside the Vertex-I and Vertex-II magnets is perpendicular to the figure. The electric drift field of the TPCs is also perpendicular to the figure. The GRC system for in-situ drift velocity monitoring was also conceptualized and introduced during the LS2. The GRC system, as of now, consists of two adjacent identical chambers (GRC-1 and GRC-2), for redundancy reasons.
@@
The concept of the Geometry Reference Chambers (GRCs), just downstream of MTPC-L, was also developed during the LS2 upgrade period. Its lightweight design was motivated by short development time, and relatively fast and low risk implementability. In addition, the TPC chambers were moved during the LS2 upgrade, and therefore seven calibration constants apart from the drift velocity must be re-determined. These are the trigger latency, the three-dimensional alignment shifts, and the three alignment angles for each TPC. These are too many unknowns to determine without an absolute reference. The GRC enables unambiguous determination of all seven parameters and the drift velocity.
@@ 3 The GRC system
The NA61/SHINE TPCs have a drift length of ${\sim}1\,\mathrm{m}$ and are capable of resolving particle cluster positions at the ${\sim}1\,\mathrm{mm}$ at the extremum of the drift volume. Therefore, in order to reduce the systematic errors of the position measurements below the position resolution, the in-situ drift velocity has to be determined with permil accuracy. The drift velocity in a typical TPC system is affected by the working gas composition, the electric drift field applied to the TPC, the gas temperature, and the gas pressure. Most large TPC systems, such as the TPCs in NA61/SHINE experiment or the one in ALICE [5] at CERN’s LHC, or in STAR [6, 7] at BNL’s RHIC, are equilibrated with the atmospheric pressure, with a very slight constant overpressure applied in order to avoid air inflitration. Since the gas composition, drift field and the temperature is typically stabilized, the shortest time scale variations are caused by meteorological changes in the ambient air pressure. This means that the time scale of anticipated substantial drift velocity variations are not shorter than about $5$ minutes. Therefore, the design goal of the GRC system was to have an in-situ drift velocity measurement inside MTPC-L with a permil statistical error for about every $5$ minute time window during data taking. Given the worst-case minimal particle flux, this requirement gives the lower bound to the acceptance (area) of the GRC in case of low-multiplicity collision types. In order to minimize the necessary area of the GRC, a location was chosen with the maximum possible particle flux passing through MTPC-L, but avoiding the vicinity of the beam spot. This justifies the chosen placement of the GRC system, shown in Fig.2. At the selected location, the worst-case minimum particle flux was estimated using formerly recorded proton-proton collision data at $13\,\mathrm{GeV}/c$ beam momentum. With about a factor of two safety margin, this resulted in a $40\,\mathrm{cm}$$\times$ $120\,\mathrm{cm}$ sensitive area, with the longer dimension fully covering the drift direction of MTPC-L. For redundancy reasons, two identical GRC stations (GRC-1 and GRC-2) were placed in an adjacent position (see Fig.2).
@@
For cost efficiency reasons, the GRC implementation was chosen to be a robust MWPC with cartesian readout loosely based on the design [8, 9], which was originally optimized for cosmic muon imaging purposes. Gas amplification is achieved on the anode wires (Sense Wires), and the transverse position information is read from adjacent wires (Field Wires) parallel to these in the same plane. For the vertical position information relevant for drift velocity monitoring, the segmentation is achieved using wires with the same functionality as cathode strips, perpendicular to the Sense Wire / Field Wire direction (those are called Pickup Wires). The design choice to use such wires instead of strips is due to easier construction of the whole chamber. For low-multiplicity data taking campaigns, the entire GRC acceptance is read out as is. For high-multiplicity runs, such as heavy-ion data taking periods, the cartesian combinatorial background becomes an issue. In order to mitigate this, the design allowed for high-voltage jumpers used to turn off individual Sense Wires. This has the effect of narrowing the transverse acceptance down to a single active sense wire. This feature completely eliminates the cartesian combinatorics, preserving the drift velocity determination accuracy in high multiplicity runs. The concept is sketched in the left panel of Fig.3.
@@
Figure 3: (Color online) Left panel: sketch of a GRC chamber as seen from the direction of the arriving particle flux (not to scale). The transverse direction is covered by the Field Wire / Sense Wire plane, where the electron multiplication takes place on the Sense Wires, and their mirror charges on the Field Wires are read out. The TPC drift direction is covered by the Pickup Wire plane, also seeing the mirror charges of the charge clouds on the Sense Wires. This provides the other cartesian coordinate for readout. Individual Sense Wires can be connected to high voltage or ground using configurable high-voltage jumpers (not drawn on the figure). Right panel: sketch of a GRC as seen from side view (not to scale). The thickness of the drift region is larger in the direction of the particle flux, and a drift cathode electrode with adjustable potential is used. This combination makes sure that the charge collection can be sufficiently slowed down such that the signal in the chamber persists until the TPC readout electronics, used for the GRC readout, begins to measure.
@@
In order to reduce cost and development time for the readount electronics, the GRCs were designed to be read out using the existing TPC FEEs. Two TPC FEE cards are enough to read out the induced charges on the $32$ Field Wire channels $\times$ $192$ Pickup Wire channels in each GRC. A natural issue with this solution is that the TPC, being a relatively slow device, allows for a relatively large trigger latency. A typical TPC readout FEE is generally optimized for making meaningful charge sampling measurements only $1-2\,\mu\mathrm{s}$ after the beam passage. In order to make sure that the signal in the GRC persists until the TPC FEE begins to take meaningful charge samples, a $30\,\mathrm{mm}$ thick drift volume was added. This has the effect of extending charge collection over time. In order to control the signal collection timescale, an adjustable drift field electrode was added to the design. The signal stretching concept is shown schematically in the right panel of Fig.3. The GRC front-end readout system is fully integrated with the NA61/SHINE Data Acquisition System (DAQ). GRC-specific event reconstruction software is fully integrated with the NA61/SHINE offline software (ShineOffilne) [10]. The fundamental parameters of the MWPC design of the GRCs are listed in Table 1.
@@
Since the GRC performs a differential measurement along the TPC drift direction, precise alignment of the GRCs with respect to the TPC is not critical. However, if the TPC-GRC alignment is known, the GRC system can be used to determine a further calibration constant: the trigger latency ($t_{0}$). In order to make use of this possibility, care was taken in the design such that the position of the optical survey target for geodesic localization of the GRCs can be accurately related to the internal wire positions. A CAD sketch of the sensor plane and a photograph of the installed GRC system is seen in Fig.4.

### v2

@@ Abstract
This paper presents a novel method for low maintenance, low ambiguity in-situ drift velocity monitoring in large volume Time Projection Chambers (TPCs). The method was developed and deployed for the $40\,\mathrm{m}^{3}$ TPC tracker system of the NA61/SHINE experiment at CERN, which has a one meter of drift length. The method relies on a low-cost multi-wire proportional chamber placed next to the TPC to be monitored, downstream with respect to the particle flux. Reconstructed tracks in the TPC are matched to hits in the monitoring chamber, called the Geometry Reference Chamber (GRC). Relative differences in positions of hits in the GRC are used to estimate the drift velocity, removing the need for an accurate alignment of the TPC to the GRC. An important design requirement on the GRC was minimal added complexity to the existing system, in particular, compatibility with Front-End Electronics cards already used to read out the TPCs. Moreover, the GRC system was designed to operate both in large and small particle fluxes. The system is capable of monitoring the evolution of the drift velocity inside the TPC down to a one permil precision, with a few minutes of data collection.
@@ 1 Introduction
The NA61 experiment, also called the SPS Heavy Ion and Neutrino Experiment (SHINE) [1], is a large-acceptance hadron spectrometer experiment at the Super Proton Synchrotron (SPS) accelerator at CERN. Its physics program includes measurements for heavy-ion physics, as well as measurements on particle production in hadron-nucleus collisions, with an emphasis on the application of those measurements in flux predictions for long-baseline neutrino oscillation experiments (T2K, DUNE, NOvA, Hyper-Kamiokande) [2, 3, 4, 5] and for cosmic ray observatories (such as Pierre Auger Observatory) [6]. These physics programs demand the operation of NA61/SHINE in both high and low particle multiplicity environments.
@@
The main components of the NA61/SHINE experiment are two large superconducting bending magnets and a system of large volume Time Projection Chambers (TPCs). As the drift dimension of the TPCs is considerably large (${\gtrsim}1\,$meter), in order to achieve necessary position resolution in the drift direction, the drift velocity inside the TPCs needs to be monitored at the permil level. Since the drift velocity slowly evolves with time, mostly due to variations in the ambient conditions or gas composition, its evolution needs to be followed every few minutes of data taking. In this paper, we describe a cost-efficient supplementary detector system specifically developed for this purpose.
@@
One of the default methods for monitoring drift velocity in a TPC is to analyze its exhaust gas with a small probe chamber [7, 1, 8, 9] and then predict the drift velocity inside the TPC via the known pressure, temperature and drift field in the TPC gas. This method dates back to the large-volume TPCs of ALEPH and DELPHI at LEP [10, 11]. NA61/SHINE also uses this technique to monitor the drift velocity in its TPCs, but due to the limited absolute systematic accuracy of the method, it is mainly used for monitoring the time stability of the gas composition. [^1] The so-called normalized drift velocity, estimated via exhaust analyzer, is the hypothetical drift velocity extrapolated to a fixed pressure, fixed temperature, and to a fixed drift field. That quantity, although is not directly related to the actual drift velocity inside the TPC, but is rather useful for monitoring the time stability of the working gas composition. The exhaust method is prone to uncertainties in pressure and temperature measurements in the large-volume chambers, and is also vulnerable to contaminations along the exhaust collection and system outgassing.
An other typical method applied in a number of experiments is the so-called bottom point method: the drift time of the last points of tagged cathode-exiting tracks can be used for drift velocity estimation, provided that the delay of the trigger signal to the TPC readout is known to a good accuracy, along with a precise estimate for the effective drift length. In NA61/SHINE we do not rely on this method, primarily because it is prone to systematic uncertainties of the effective drift length, but more importantly of the trigger delay, which is not even known a priori. A further method, often applied in large volume TPC experiments, relies on imaging start and end points of cosmic ray tracks. In NA61/SHINE, such method is not suitable, as the orientation of the TPC acceptance and drift direction is suboptimal with respect to the celestial directions, leading to large uncertainties in the cosmic track start and end point reconstruction.
@@
Due to the systematic uncertainties of the above approaches, a more advanced absolute in-situ drift velocity measurement method is needed. For such purpose, typically a UV laser system is used: ionizing laser beam tracks are projected into the sensitive volume at known locations, and the drift velocity can then be measured from the apparent positions of these beams in terms of TPC drift time. Direct measurement methods using laser systems are common [12, 8] and very accurate. UV laser-based drift velocity and geometry monitoring systems, however, are rather hard to implement or upgrade on existing chambers retrospectively. Moreover, the cost and development time for such systems is substantial. NA61/SHINE is not equipped with such a UV laser based system.
@@
Along with other CERN experiments, the NA61/SHINE facility was also significantly upgraded during the Long Shutdown 2 (LS2) period (2019-2021) of the accelerator complex. Since during the LS2 the subdetectors were moved, precise alignment calibration of the chambers was required, as was a system for accurately and reliably determining their drift velocities. As the TPCs were already constructed, minimally-invasive measurement systems were preferred. Our choice was a simple solution: a planar detector with fixed segmentation along the TPC drift direction, called the Geometry Reference Chamber (GRC), was to be placed downstream of the TPC system with respect to the particle flux. TPC tracks reconstructed with an approximately known drift velocity are extrapolated to the GRC and paired with the GRC hits. If the drift velocity estimate differs from the true drift velocity inside the TPC, the mismatch of the drift coordinate of the hits becomes worse with increasing drift depth. The ratio between the initial approximate drift velocity and the actual drift velocity in the chamber can then be estimated from the slope of this mismatch scatter plot. That is, the GRC is used as a differential length scale in order to convert the drift time, measured by the TPC readout, to drift depth. It is quite advantageous that the GRC method is differential: an accurate alignment of the TPC against the GRC is not needed. The concept is illustrated in Fig.1.222Once the drift velocity has been calibrated with such method in one of the TPC chambers in the detector complex, the other chambers can be calibrated against each-other, using the already calibrated chambers as geometry reference. Moreover, the remaining geometry calibration constants of the TPC system can also be estimated, such as the trigger delay ($t_{0}$), alignment shift and alignment angles for each chamber, see Appendix A and B.
@@
Figure 1: (Color online) Illustration of the Geometry Reference Chamber (GRC) concept. The particle tracks in the TPC chamber are reconstructed with an initial estimate for the drift velocity and are paired to the hits of a segmented detector, the GRC. In case the initial approximate drift velocity estimate differs from the true drift velocity inside the TPC, the mismatch along the drift coordinate gets gradually worse at increasing drift depths. The slope of the drift coordinate mismatch against the drift coordinate as measured by the GRC provides the estimate for the corresponding multiplicative correction factor. (The figure shows the projection orthogonal to the magnetic bending plane, in case the TPC is used in a magnetic spectrometer experiment, such as the NA61/SHINE.)
@@
In order to implement the GRC concept, several design requirements were considered. To reduce the development time and cost of the readout electronics, a solution using the existing TPC readout Front-End Electronics (FEE) was chosen. To reduce the number of readout channels, we decided to implement a simple planar Multi-Wire Proportional Chamber (MWPC) with Cartesian readout. This setting required two issues to be addressed. Firstly, the signal formation in the GRC had to be slowed down sufficiently such that the signal persists until the TPC electronics begins to sample the charges. This is not entirely trivial to achieve, since TPC readout electronics typically allow for a microsecond-level trigger delay with respect to the particle passage time, during which the signal may already disappear from the GRC. In order to mitigate this issue, a drift volume with adjustable drift field was introduced in the GRC design, transverse to the particle passage direction, and with that the GRC signal duration could be controlled. Secondly, the Cartesian readout scheme tends to be problematic whenever the particle occupancy is large. In order to overcome this limitation, the GRC acceptance was designed to be variable: for low-multiplicity collisions, the entire GRC acceptance can be used, whereas for high-multiplicity collisions, individual anode wires (sense wires) of the GRCs can be disabled, down to a single amplifying wire. This helps to reduce the number of combinatorial ghost hits.
@@
The structure of the paper is as follows. In Section 2 an overview of the NA61/SHINE experimental facility is given. In Section 3 the GRC design requirements, parameters, and implementation are discussed. In Section 4 the operational performance is shown. In Section 5 a conclusion is provided. Appendices A and B discuss some details of systematic errors and the estimation of further geometry-related TPC calibration constants, where the GRC chamber plays a role.
@@ 2 The NA61/SHINE experimental facility
The NA61/SHINE [1] is a fixed-target hadron spectrometer experiment at the CERN SPS accelerator. Large parts of its tracking devices were inherited from a previous experiment called NA49 [8]. Its physics program covers the study of strongly-interacting matter via heavy-ion collisions, and measurements of identified particle production spectra in hadron-nucleus collisions as reference data for flux prediction in long baseline neutrino experiments and large area cosmic ray observatories.
@@
An overview of the NA61/SHINE experiment is shown in Fig.2. Two large superconducting bending magnets (Vertex-I and II) are responsible for particle deflection for charge and momentum determination. Their total maximum bending power is ${\sim}9\,\mathrm{Tm}$ (up to $1.5\,\mathrm{Tesla}$ magnetic field in Vertex-I and $1.1\,\mathrm{Tesla}$ in Vertex-II). A target holder with target-in / target-out moving capability sits just upstream of the first Vertex TPC, upstream/downstream always meant with respect to the beam direction throughout the paper. Thin targets can be placed inside a silicon Vertex Detector (VD) for precise vertex determination. NA61/SHINE also has the ability to measure interactions in extended replica targets for long baseline neutrino experiments. The tracking devices for spectrometry are composed of eight large volume TPCs (total ${\sim}40\,\mathrm{m}^{3}$ and ${\sim}1\,\mathrm{m}$ drift length), capable of performing both tracking and $\mathrm{d}E/\mathrm{d}x$ measurements. A Multigap Resistive Plate Chamber (MRPC) based Time-of-Flight wall (L-ToF) provides further particle identification (PID) capabilities around mid-rapidity, whereas a scintillator based Time-of-Flight wall (F-ToF) covers the forward phase space, enabling two-dimensional (ToF+$\mathrm{d}E/\mathrm{d}x$) separation of particle species at large parts of the acceptance. A calorimeter is placed at the end of the beamline, called the Projectile Spectator Detector (PSD), which helps to characterize collision centrality in heavy-ion collisions, and consists of two compartments (Main- and Forward-PSD, also called MPSD and FPSD) for optimal shower containment. Upstream of the target position, a set of beam position detectors (BPDs) provide beam tracking information, while scintillator and Cherenkov detectors serve as beam trigger with PID capability (not shown in the figure). On the beamline between VTPC-1 and the GapTPC, a small plastic scintillator with a $1\,\mathrm{cm}$ diameter serves as an interaction trigger in most collision types (S4). In rare run settings, a very similar scintillator (S5) just upstream of MPSD is used instead. For heavy-ion runs, the scintillator (S3) just downstream of VD is used for interaction definition through detecting loss of charge (i.e., by sensing a decreased $Z^{2}$ with respect to the beam nuclei). The VD, the new silicon BPDs, the MRPC-based L-ToF, the MPSD and FPSD were introduced during the LS2 upgrade period between 2019-2021. Also during LS2, the data acquisition (DAQ) was upgraded, allowing for an event recording rate up to $1.8\,\mathrm{kHz}$ for the detector complex.
@@
Figure 2: (Color online) The NA61/SHINE detector configuration, after the Long Shutdown 2 (LS2) upgrade, i.e. since 2022. The drawing shows the top view of the bending plane. That is, the approximately homogeneous magnetic field inside the Vertex-I and Vertex-II magnets is perpendicular to the figure (${\sim}y$ axis). The electric drift field of the TPCs is also perpendicular to the figure (${\sim}y$ axis). The GRC system for in-situ drift velocity monitoring was also conceptualized and introduced during the LS2. The GRC system, as of now, consists of two adjacent identical chambers (GRC-1 and GRC-2), for redundancy reasons. As the figure is a top view of the experiment, the shorter dimension of the GRCs is seen in this outline.
@@
The concept of the Geometry Reference Chambers (GRCs), just downstream of MTPC-L, was also developed during the LS2 upgrade period. Its lightweight design was motivated by short development time, and relatively fast and low risk implementability. In addition, the TPCs were moved during the LS2 upgrade, and therefore seven calibration constants apart from the drift velocity must be re-determined for each TPC. These are the trigger delay, the three-dimensional alignment shifts, and the three alignment angles, per chamber, independently. The large number of unknown calibration constants are hard to determine without an absolute reference. The GRC enables unambiguous determination of all seven parameters and the drift velocity, for each component of the TPC system.
@@ 3 The GRC system
The NA61/SHINE TPCs have a drift length of ${\sim}1\,\mathrm{m}$ and are capable of resolving particle cluster positions at the ${\sim}1\,\mathrm{mm}$ at the extremum of the drift volume. Therefore, in order to reduce the systematic errors of the position measurements below the position resolution, the drift velocity has to be determined with permil accuracy. The drift velocity in a typical TPC system is affected by the working gas composition, the electric drift field applied to the TPC, the gas temperature, and the gas pressure. Most large TPC systems, such as the TPCs in NA61/SHINE experiment or the one in ALICE [12] at CERN’s LHC, or in STAR [13, 14] at BNL’s RHIC, are equilibrated with the atmospheric pressure, with a very slight constant overpressure applied in order to avoid air infiltration. Since the gas composition, drift field and the temperature is typically stabilized, the shortest time scale variations are caused by meteorological changes in the ambient air pressure. This means that the time scale of anticipated substantial drift velocity variations are not shorter than about $5$ minutes. Therefore, the design goal of the GRC system was to have a drift velocity measurement inside MTPC-L with a permil statistical error for about every $5$ minute time window during data taking. Given the worst-case minimal particle flux, this requirement gives the lower bound to the acceptance (area) of the GRC in case of low-multiplicity collision types. In order to minimize the necessary area of the GRC, a location was chosen with the maximum possible particle flux passing through MTPC-L, but avoiding the vicinity of the beam spot. This justifies the chosen placement of the GRC system, shown in Fig.2. At the selected location, the worst-case minimum particle flux was estimated using formerly recorded proton-proton collision data at $13\,\mathrm{GeV}/c$ beam momentum. With about a factor of two safety margin, this resulted in a $40\,\mathrm{cm}$$\times$ $120\,\mathrm{cm}$ sensitive area, with the longer dimension fully covering the drift direction of MTPC-L. For redundancy reasons, two identical GRC stations (GRC-1 and GRC-2) were placed in an adjacent position (see Fig.2).
@@
For cost efficiency reasons, the GRC implementation was chosen to be a robust MWPC with Cartesian readout loosely based on the design [15, 16], which was originally optimized for cosmic muon imaging purposes. Gas amplification is achieved on the anode wires (sense wires), and the transverse position information is read from adjacent wires (field wires) parallel to these in the same plane. For the vertical position information relevant for drift velocity monitoring, the segmentation is achieved using wires with the same functionality as cathode strips, perpendicular to the sense wire / field wire direction (those are called pickup wires). The design choice to use such wires instead of strips is due to easier construction of the whole chamber. For low-multiplicity data taking campaigns, the entire GRC acceptance is read out as is. For high-multiplicity runs, such as heavy-ion data taking periods, the Cartesian combinatorial background becomes an issue. In order to mitigate this, the design allows for high-voltage jumpers used to turn off individual sense wires: the sense wires can be connected either to their operational voltage or to the chamber ground. The jumpers are configured manually at the start of a data taking campaign. Turning off sense wires has the effect of narrowing the transverse acceptance, eventually down to a single active sense wire. This feature completely eliminates the Cartesian combinatorics, preserving the drift velocity determination accuracy in high multiplicity runs. The concept is sketched in the left panel of Fig.3.
@@
Figure 3: Left panel: sketch of a GRC chamber as seen from the direction of the arriving particle flux (not to scale). The transverse direction is covered by the field wire / sense wire plane, where the electron multiplication takes place on the sense wires, and their mirror charges on the field wires are read out. The TPC drift direction is covered by the pickup wire plane, also seeing the mirror charges of the charge clouds on the sense wires. This provides the other Cartesian coordinate for readout. Individual sense wires can be connected to high voltage or ground using configurable high-voltage jumpers (not drawn on the figure). Right panel: sketch of a GRC as seen from side view (not to scale). The thickness of the drift region is larger in the direction of the particle flux, and a drift cathode electrode with adjustable potential is used. This combination makes sure that the charge collection can be sufficiently slowed down such that the signal in the chamber persists until the TPC readout electronics, used for the GRC readout, begins to measure.
@@
In order to reduce cost and development time for the readout electronics, the GRCs were designed to be read out using the existing TPC FEEs. Two TPC FEE cards are enough to read out the induced charges on the $32$ field wire channels $\times$ $192$ pickup wire channels in each GRC. A natural issue with this solution is that the TPC, being a relatively slow device, allows for a relatively large trigger delay. Typical TPC trigger and readout electronics are only optimized for making meaningful charge sampling measurements $1-2\,\mu\mathrm{s}$ after the beam passage. This relatively large but constant time delay is due to trigger logic decision time, signal arrival time on the signal cables, as well as the startup time of the FEE logic. In order to make sure that the signal in the GRC persists until the TPC FEE begins to take meaningful charge samples, a $30\,\mathrm{mm}$ thick drift volume was added. This has the effect of extending charge collection over time. In order to control the signal collection timescale, an adjustable drift field electrode was added to the design. The signal stretching concept is shown schematically in the right panel of Fig.3.333Due to the mentioned thickened charge collection volume, the charge collection direction slightly differs from the the particle incidence direction in case of tracks arriving non-orthogonally at the GRC surface. This causes a small angle-dependent position bias, which can be easily corrected for, using the incidence angles estimated by the TPC track piece. Due to the NA61/SHINE geometry, the pertinent bias is of the order of ${\sim}1\,\mathrm{mm}$ at the extremes, and vanishes toward the middle of the GRC acceptance. Moreover, in the NA61/SHINE geometry, this bias is second order, therefore it happens to cancel in the GRC-based drift velocity calibration procedure. The GRC front-end readout system is fully integrated with the NA61/SHINE Data Acquisition System (DAQ). GRC-specific event reconstruction software is fully integrated with the NA61/SHINE offline software (ShineOffilne) [17]. The fundamental parameters of the MWPC design of the GRCs are listed in Table 1.
@@
Since the GRC performs a differential measurement along the TPC drift direction, precise alignment of the GRCs with respect to the TPC is not critical. However, if the TPC-GRC alignment is known, the GRC system can be used to determine a further calibration constant: the trigger delay ($t_{0}$). In order to make use of this possibility, care was taken in the design that the position of the optical survey target for geodesic localization of the GRCs can be accurately related to the internal wire positions. A CAD sketch of the sensor plane and a photograph of the installed GRC system is seen in Fig.4.